For a certain gas, R = 0.277 kJ/kg-K and k = 1.384. a) What are the value of cp

and cv? b) What mass of this gas will occupy a volume of 0.425 m3 at 517.11

kPaa and 26.7 0C? c) If 31.65 kJ are transferred to this gas at constant

volume, what are the resulting temperature and pressure?


1
Expert's answer
2022-02-14T14:23:38-0500

We know that


"n=\frac{PV}{RT}=\frac{517.11\times0.425}{0.277\times10^3\times299.7}=2.64\times10^{-3}mol"

Idial gas equation

"Q=nC_v\u2206T"

"C_v=\frac{Q}{n∆T}"

"C_v=\frac{31.65}{2.64\times10^{-3}\times299.7}=40.002"

Gas relationship

"C_p-C_v=R"

"C_p=C_v+R"


"C_p=40.002+0.277=40.279KJ\/kg-K"

(B)

Mass


"m=\frac{PV}{RT}=\frac{517.11\times10^{3}\times0.425}{0.277\times10^3\times200.7}=3.95kg"

(C)


"V=\frac{nRT}{P}=\frac{2.64\times10^{-3}\times0.277\times10^{3}\times299.7}{517.11\times10^{3}}=4.23\times10^{-4}m^3"